OpenStudy (anonymous):

Solve 5x2 = -30x - 65. x = -3 ± 2i x = -3 ± 4i x = -6 ± 2i x = -6 ± 4i

OpenStudy (anonymous):

@danish071996

OpenStudy (anonymous):

its 5x^2 or 5x2?

OpenStudy (anonymous):

@danish071996 5x^2

OpenStudy (anonymous):

Solve x2 + 4x + 8 = 0. x equals negative 2 plus or minus 2 I x equals negative 2 plus or minus 4 I x equals negative 4 plus or minus 2 I x equals negative 4 plus or minus 4 I

OpenStudy (anonymous):

its 41 ?in last option

OpenStudy (anonymous):

Solve -3x2 + 30x - 90 = 0. x equals 5 plus or minus 2i square root of 5 x equals 5 plus or minus i square root of 5 x equals 10 plus or minus 2i square root of 5 x equals 10 plus or minus i square root of 5

OpenStudy (anonymous):

Solve x2 - 3x = -8. x equals quantity of 3 plus or minus I square root of 29 all over 2 x equals quantity of 3 plus or minus I square root of 23 all over 2 x equals quantity of negative 3 plus or minus I square root of 29 all over 2 x equals quantity of negative 3 plus or minus I square root of 23 all over 2

OpenStudy (anonymous):

3x2^2 and x^2

OpenStudy (anonymous):

@danish071996

OpenStudy (michelleceirra606):

x = -2.5